OpenStudy (anonymous):

Find the volume of the solid obtained by rotating the region bounded by the given curve about the y-axis. y=3xe^-x y=0 x=3 about the y=axis I was able to get the graph and all I really need is someone to explain how to set up the integral. Thanks for your help!

OpenStudy (john_es):

Do yo mean, \[y=3xe^{-x}\] or \[y=3e^{-x}\]?

OpenStudy (anonymous):

the first one

OpenStudy (john_es):

with one dimensional integration?

OpenStudy (anonymous):

yes

OpenStudy (john_es):

Sorry, I don't see an easy way to solve this in one dimension.

OpenStudy (john_es):

Around the y axis. Around the x axis it should be easy.

OpenStudy (anonymous):

the question says around the y-axis is it easier as a 3D figure?

OpenStudy (john_es):

As a 3D figure, integrating in three dimensions, I think it should be easier. The problem I see when you integrate around y axis is that you need the funtion f(y), but the equation you give don't let us put x in function of y, so we cannot find an easy expression for f(y). Is this problem from a book or could you tell me from what course is?

OpenStudy (anonymous):

its from calc 2

OpenStudy (john_es):

What textbook do you use?

OpenStudy (anonymous):

Calculus Early Transcendentals 7E by james stewart

OpenStudy (john_es):

Ok, I will check it. May be we'll have some luck.

OpenStudy (john_es):

Could you please indicate me the chapter?

OpenStudy (anonymous):

6.3

OpenStudy (john_es):

Ok, I completely forget about it. But it is easy. It is integration using cylindrical shells. The volume generated by a function f(x) (like the one they give you, where you cannot put x as a function of y), is, \[V=\int_a^b2\pi x f(x)dx\]So in this case, \[V=\int_0^32\pi x (3xe^{-x})dx=6\pi\int_0^3 x^2e^{-x}dx\] Could you integrate it?

OpenStudy (anonymous):

i think so thanks! I was just really unsure of the e

OpenStudy (john_es):

Ok, only to check it I obtain the value, and find, \[V=6 \left(2-\frac{17}{e^3}\right) \pi\] I hope it helps you.

OpenStudy (anonymous):

it did! Could you help me with one other problem?

OpenStudy (john_es):

I'll try it :)

OpenStudy (anonymous):

The integral represents the volume of a solid. Describe the solid. \[2\pi \int\limits_{0}^{6} \frac{ 7y }{ 7+y^2 } dy\] The solid is obtained by rotating the region \[o \le x \le ? \] \[0 \le y \le 6 \] about the x-axis using cylindrical shells.

OpenStudy (anonymous):

I was able to figure out the y and the x-axis part but i don't see how to sole for the x part

OpenStudy (john_es):

To solve for the x part we need the function f(x). If you compare the integral you give with the one we use to integrate volumes around the y axis, \[V=\pi\int f(y)^2dy\]We know that \[x=f(y)\]In our case, \[f(y)^2=2\frac{7y}{7+y^2}=\frac{14y}{7+y^2}\Rightarrow f(y)=\sqrt{\frac{14y}{7+y^2}}\]So, \[x=\sqrt{\frac{14y}{7+y^2}}\]Now you should only substitute the y with value 6, and you'll obtain the required result.

OpenStudy (john_es):

My result is, to check, \[2 \sqrt{\frac{21}{43}}\]

OpenStudy (anonymous):

It says its wrong

OpenStudy (anonymous):

\[0 \le x \le 2\sqrt{\frac{ 21 }{ 43 }}\] is not right

OpenStudy (anonymous):

wait shouldnt V= 2pi not just pi

OpenStudy (john_es):

May be they don't simplify, \[\sqrt{84/43}\]

OpenStudy (john_es):

Does your integral has a 2 or not? :)

OpenStudy (anonymous):

\[V=2\pi \int\limits_{a}^{b}y f(y) dy \]

OpenStudy (anonymous):

thats the equation the question gives

OpenStudy (john_es):

Ah, and the f(y) you posted before?

OpenStudy (john_es):

Anyways, if the problem gives you the last expression, could you write f(y)?

OpenStudy (anonymous):

yea the integral from 0 to 6

OpenStudy (anonymous):

i thought it was \[\sqrt{\frac{ 7 }{ x-7 }}\] but it said it was wrong

OpenStudy (john_es):

Ok, I think it should be this way, \[V=2\pi\int_0^6yf(y)dy\Rightarrow x=\frac{7}{7+y^2}\Rightarrow x=\frac{7}{7+36}=7/43\]

OpenStudy (john_es):

Ok, I think it should be this way, \[V=2\pi\int_0^6yf(y)dy\Rightarrow x=\frac{7}{7+y^2}\Rightarrow x=\frac{7}{7+36}=7/43\]

OpenStudy (anonymous):

got it wrong the answer is actually 7/7+y^2 not suppose to plug in the 6 but thanks for your help!!

OpenStudy (john_es):

Oh, of course. I was thinking in a numeric limit! But they ask you for the function! Sorry. ;)

OpenStudy (anonymous):

its fine thanks for your help :) I only have one more problem to figure out so if i get this one i think ill be ok lol

OpenStudy (john_es):

Well, post it and we'll try :)

OpenStudy (anonymous):

The region bounded by the given curves is rotated about the specified axis. Find the volume V of the resulting solid by any method. \[y=-x^2+8x-15 \] y=0 about the x-axis

OpenStudy (john_es):

What is the other curve and where are the limits of integration?

OpenStudy (anonymous):

thats all the question gives :/

OpenStudy (anonymous):

since y=0 the intersection points are (0,3) and (0,5)

OpenStudy (john_es):

True. I just read curves and was thinking in another parabola XD.

OpenStudy (john_es):

Then it should be easy, \[V=\pi\int_3^5 (-x^2+8x-15)^2dx\]

OpenStudy (anonymous):

that was so simple i guess i was over thinking it haha

OpenStudy (john_es):

Sometimes happens ;)